OpenStudy (darkbluechocobo):

what kind of number is 4+sqrt-25

jimthompson5910 (jim_thompson5910):

What happens when you try to evaluate \(\Large \sqrt{-25}\) ?

OpenStudy (darkbluechocobo):

you get -5i

OpenStudy (triciaal):

and therefore complex number

OpenStudy (darkbluechocobo):

because it is a mixture between imaginary and real number?

OpenStudy (triciaal):

good way to put it yes

jimthompson5910 (jim_thompson5910):

not -5i, just 5i

jimthompson5910 (jim_thompson5910):

\(\Large \sqrt{-25} = 5i\) \(\Large -\sqrt{-25} = -5i\)

OpenStudy (triciaal):

general form for complex number a + bI

OpenStudy (darkbluechocobo):

oh sorry i accidently put that there. I dont know why i put a -5

OpenStudy (triciaal):

a and b are real I = rt of -1

OpenStudy (triciaal):

@jim_thompson5910 I think both would be correct rt-25 = rt-1* rt25 using +ve and neg values for rt 25 -5I and 5I

jimthompson5910 (jim_thompson5910):

no, the square root function only produces one output for any given input I'm using the complex number version of the square root function

OpenStudy (triciaal):

are you saying there is a rule when using complex numbers to use one or the other ?

OpenStudy (triciaal):

if using a calculator will give the absolute value yes (-x)^2 = (x)^2

jimthompson5910 (jim_thompson5910):

it is true that (5i)^2 = -25 and (-5i)^2 = -25 therefore, x^2 = -25 has the solutions x = 5i or x = -5i

jimthompson5910 (jim_thompson5910):

however, the square root function only returns one exact value when you input a number example: sqrt(9) = 3 it's not sqrt(9) = 3 or -3

jimthompson5910 (jim_thompson5910):

if you want the plus or minus, you have to explicitly tack it on \(\Large \pm\sqrt{9} = \pm 3\)